# Decomposing Disease Descriptions for Enhanced Pathology Detection: A Multi-Aspect Vision-Language Pre-training Framework

### v3

@@ Abstract
Medical vision language pre-training (VLP) has emerged as a frontier of research, enabling zero-shot pathological recognition by comparing the query image with the textual descriptions for each disease. Due to the complex semantics of biomedical texts, current methods struggle to align medical images with key pathological findings in unstructured reports. This leads to the misalignment with the target disease’s textual representation. In this paper, we introduce a novel VLP framework designed to dissect disease descriptions into their fundamental aspects, leveraging prior knowledge about the visual manifestations of pathologies. This is achieved by consulting a large language model and medical experts. Integrating a Transformer module, our approach aligns an input image with the diverse elements of a disease, generating aspect-centric image representations. By consolidating the matches from each aspect, we improve the compatibility between an image and its associated disease. Additionally, capitalizing on the aspect-oriented representations, we present a dual-head Transformer tailored to process known and unknown diseases, optimizing the comprehensive detection efficacy. Conducting experiments on seven downstream datasets, ours improves the accuracy of recent methods by up to 8.56% and 17.0% for seen and unseen categories, respectively. Our code is released at https://github.com/HieuPhan33/MAVL.
@@ 3.1 Problem setting
where $t_{i}$ refers to the number of entities in the report, and $\text{location}_{t}$ is the location (e.g. lung, spine) where the medical $\text{entity}_{t}$ occurs.
@@ 5.1 Zero-shot evaluation
Classification for unseen diseases. Table 2 presents the zero-shot classification benchmark of unseen diseases. We include the performance of both supervised and contrastive heads in our model for comparisons. We note three observations. First, both our heads consistently outperform previous methods on both datasets. Our supervised head drastically improves the AUC of a contrastive CheXzero, and a supervised MedKLIP by up to 7.66% and 12.49%. By decomposing the unstructured disease description into a common set of structured elemental aspects, our model effectively links new disease aspects with base diseases, improving feature representations of unseen diseases.
Second, contrastive models are generally more effective for novel disease recognition than the supervised counterparts. While the supervised MedKLIP struggles to classify many unseen diseases in PadChest, our contrastive head improves the AUC score of MedKLIP and CheXzero by 16.76% and 5.58%, respectively. This shows the effectiveness of our dual-head Transformer design, tailored to process both known and unknown diseases. Lastly, both of our heads surpasses previous methods in classifying rare diseases, significantly improving the AUC scores of MedKLIP and CheXzero by 17.0% and 7.42%. Leveraging prior knowledge about pathological visual profiles boosts the classification of the long-tailed distribution.
@@ 5.2 Fine-tuning evaluation
This section provides fine-tuning evaluations by using the pre-trained models of different VLP methods to initialize weights and fine-tune downstream datasets. Table 4 presents fine-tuning results on 4 datasets with 1%/10%/100% data portion, which is consistent with previous works [39, 6]. Our model yields consistent improvements over previous works across all settings, especially under 1% data portion for both classification and segmentation. Notably, the proposed MAVL significantly outperforms the second-best method, MedKLIP, by 7.38% and 12.82% on SIIM Pneumothorax classification and RSNA-Pneumonia segmentation when fine-tuning on 1% data. On a fine-grained ChestXray-14 with 14 lung diseases, our MAVL offers drastic performance gain even when fine-tuning with 100% data. Injecting prior knowledge about the disease’s visual aspects during pre-training improves the model’s transferability under low-shot learning, especially on fine-grained disease recognition.
@@ 5.3 Ablation study
Effectiveness of multi-aspect decomposition for VLP. This section answers two questions: (i) whether adding more concepts improves zero-shot performance, and (ii) which concept matching yields higher performance gain. Fig. 6(a) presents the performance when adding a single aspect. Matching with any single aspect boosts the accuracy of matching with a disease’s definition as in [39]. Notably, the gains on on unseen PadChest and Covid CXR-2 are 3.64% and 4.02%, respectively when using the aspect ‘texture’. More distinct aspects such as texture, patterns, and shapes yield higher gains. Fig. 6(b) shows the AUC scores when adding increasing numbers of visual aspects. Matching with higher numbers of aspects improves the zero-shot classification.
Figure 6: AUC scores of our method when (a) adding a single aspect, besides a disease description and (b) when gradually adding higher numbers of aspects by their increasing importance order. The scores of using clinical description (desc.) of diseases as in [39] are included as the baseline.
Effectiveness of a dual-head design. This section analyzes the effectiveness of our dual-head network, which decouples contrastive and supervised learning. We implement a single-head model using only either a contrastive branch as in [34, 3] or a supervised branch as in [39] First, Table 5 shows that using a single-head network design as in current literature compromises the performance of either seen and unseen classes. Capitalizing both branches in our dual-head design offers the highest F1 scores of 81.68% and 26.25% respectively on unseen Covid-19 and seen ChestXray-14. Second, adding supervised learning in our Dual-Con boosts the AUC scores of the single contrastive head by 18.44% on ChestXray-14. Similarly, adding CL in our Dual-Sup improves single supervised network. Integrating both learning signals as in our dual-head design complements the learning on both seen and unseen diseases.
@@ 6 Conclusion
This study introduces a Multi-Aspect Vision-Language Matching (MAVL) framework, designed to intricately dissect pathological entities into distinct, visually descriptive aspects, thereby significantly enhancing pathology detection. The MAVL framework incorporates a dual-head, aspect-oriented Transformer, which decouples learning signals to optimize performance across both seen and unseen diseases. This approach is particularly effective for novel diseases, achieving a remarkable 16.76% performance gain over existing methods. Moreover, MAVL’s aspect-based matching leads to an impressive 17.0% AUC improvement for rare unseen diseases.

### v4

@@ Abstract
Medical vision language pre-training (VLP) has emerged as a frontier of research, enabling zero-shot pathological recognition by comparing the query image with the textual descriptions for each disease. Due to the complex semantics of biomedical texts, current methods struggle to align medical images with key pathological findings in unstructured reports. This leads to the misalignment with the target disease’s textual representation. In this paper, we introduce a novel VLP framework designed to dissect disease descriptions into their fundamental aspects, leveraging prior knowledge about the visual manifestations of pathologies. This is achieved by consulting a large language model and medical experts. Integrating a Transformer module, our approach aligns an input image with the diverse elements of a disease, generating aspect-centric image representations. By consolidating the matches from each aspect, we improve the compatibility between an image and its associated disease. Additionally, capitalizing on the aspect-oriented representations, we present a dual-head Transformer tailored to process known and unknown diseases, optimizing the comprehensive detection efficacy. Conducting experiments on seven downstream datasets, ours improves the accuracy of recent methods by up to 8.56% and 17.26% for seen and unseen categories, respectively. Our code is released at https://github.com/HieuPhan33/MAVL.
@@ 3.1 Problem setting
where $t_{i}$ refers to the number of medical entities in the report, and $\text{location}_{t}$ is the location (e.g. lung, spine) where the $\text{entity}_{t}$ occurs in the image $\mathcal{I}_{i}$.
@@ 5.1 Zero-shot evaluation
Classification for unseen diseases. Table 2 presents the zero-shot classification benchmark of unseen diseases. We include the performance of both supervised and contrastive heads in our model for comparisons. We note three observations. First, both our heads consistently outperform previous methods on both datasets. Our supervised head drastically improves the AUC of a contrastive CheXzero, and a supervised MedKLIP by up to 7.67% and 12.49%. By decomposing the unstructured disease description into a common set of structured elemental aspects, our model effectively links new disease aspects with base diseases, improving feature representations of unseen diseases.
Second, contrastive models are generally more effective for novel disease recognition than the supervised counterparts. While the supervised MedKLIP struggles to classify many unseen diseases in PadChest, our contrastive head improves the AUC score of MedKLIP and CheXzero by 16.76% and 5.58%, respectively. This shows the effectiveness of our dual-head Transformer design, tailored to process both known and unknown diseases. Third, both of our heads surpass previous methods in classifying rare diseases. Our contrastive head significantly improves the AUC scores of MedKLIP and CheXzero by 17.26% and 7.65%. Leveraging prior knowledge about pathological visual profiles boosts the classification of the long-tailed distribution.
@@ 5.2 Fine-tuning evaluation
This section provides fine-tuning evaluations by using the pre-trained models of different VLP methods to initialize weights and fine-tune downstream datasets. Table 4 presents fine-tuning results on 4 datasets with 1%/10%/100% data portion, which is consistent with previous works [39, 6]. Our model yields consistent improvements over previous works across all settings, especially under 1% data portion for both classification and segmentation. Notably, the proposed MAVL significantly outperforms the second-best method, MedKLIP, by 7.38% and 15.38% on SIIM Pneumothorax classification and RSNA-Pneumonia segmentation when fine-tuning on 1% data. On a fine-grained ChestXray-14 with 14 lung diseases, our MAVL offers drastic performance gain even when fine-tuning with 100% data. Injecting prior knowledge about the disease’s visual aspects during pre-training improves the model’s transferability under low-shot learning, especially on fine-grained disease recognition.
@@ 5.3 Ablation study
Effectiveness of multi-aspect decomposition for VLP. This section answers two questions: (i) whether adding more concepts improves zero-shot performance, and (ii) which concept matching yields higher performance gain. Fig. 6(a) presents the performance when adding a single aspect. Matching with any single aspect boosts the accuracy of matching with a disease’s definition as in [39]. Notably, the AUC score gain on unseen Covid CXR-2 is 4.02% when using the aspect ‘texture’. More distinct aspects such as texture, patterns, and shapes yield higher gains. Fig. 6(b) shows the AUC scores when adding increasing numbers of visual aspects. Matching with higher numbers of aspects improves the zero-shot classification.
Figure 6: AUC scores of our method when (a) adding a single aspect, besides a disease clinical description and (b) when gradually adding higher numbers of aspects by their increasing importance order. The scores of using clinical description (desc.) of diseases are included as the baseline.
Effectiveness of a dual-head design. This section analyzes the effectiveness of our dual-head network, which decouples contrastive and supervised learning. We implement a single-head model using only either a contrastive branch as in [34, 3] or a supervised branch as in [39] First, Table 5 shows that using a single-head network design as in current literature compromises the performance of either seen and unseen classes. Capitalizing both branches in our dual-head design offers the highest F1 scores of 81.73% and 26.25% respectively on unseen Covid-19 and seen ChestXray-14. Second, adding supervised learning in our Dual-Con boosts the AUC scores of the single contrastive head by 14.63% on ChestXray-14. Similarly, adding CL in our Dual-Sup improves single supervised network. Integrating both learning signals as in our dual-head design complements the learning on both seen and unseen diseases.
@@ 6 Conclusion
This study introduces a Multi-Aspect Vision-Language Pre-training (MAVL) framework, designed to intricately dissect pathological entities into distinct, visually descriptive aspects, thereby significantly enhancing pathology detection. Our MAVL framework incorporates a dual-head, aspect-oriented Transformer, which decouples learning signals to optimize performance across both seen and unseen diseases. This approach is particularly effective for novel diseases, achieving a remarkable 16.76% performance gain over existing methods. Moreover, MAVL’s aspect-based matching leads to an impressive 17.26% AUC improvement for rare unseen diseases.